# UMI-on-Air: Embodiment-Aware Guidance for Embodiment-Agnostic Visuomotor Policies

### v2

@@ Abstract
We introduce UMI-on-Air, a framework for embodiment-aware deployment of embodiment-agnostic manipulation policies. Our approach leverages diverse, unconstrained human demonstrations collected with a handheld gripper (UMI) to train generalizable visuomotor policies. A central challenge in transferring these policies to constrained robotic embodiments—such as aerial manipulators—is the mismatch in control and robot dynamics, which often leads to out-of-distribution behaviors and poor execution. To address this, we propose Embodiment-Aware Diffusion Policy (EADP), which couples a high-level UMI policy with a low-level embodiment-specific controller at inference time. By integrating gradient feedback from the controller’s tracking cost into the diffusion sampling process, our method steers trajectory generation towards dynamically feasible modes tailored to the deployment embodiment. This enables plug-and-play, embodiment-aware trajectory adaptation at test time. We validate our approach on multiple long-horizon and high-precision aerial manipulation tasks, showing improved success rates, efficiency, and robustness under disturbances compared to unguided diffusion baselines. Finally, we demonstrate deployment in previously unseen environments, using UMI demonstrations collected in the wild, highlighting a practical pathway for scaling generalizable manipulation skills across diverse—and even highly constrained—embodiments. All code, data, and checkpoints will be publicly released after acceptance. Result videos can be found at umi-on-air.github.io.
@@ I Introduction
There is a growing interest in extending manipulation beyond the lab and into more complex, dynamic settings. Among emerging embodiments, unmanned aerial manipulators (UAMs) hold particular promise. Essentially a manipulator with practically limitless reach, UAMs can access environments that are otherwise unreachable or unsafe—such as performing infrastructure maintenance atop towers or harvesting crops in cluttered orchards. Multiple research works have focused on those practical tasks, including non-destructive testing [bodie2019omnidirectional], painting [vempati2018paintcopter], drilling [ding2021design], light bulb installation [he2025flying], etc. These applications highlight the potential of UAMs as an embodiment, but scalable visuomotor policy learning for UAMs remains limited.
A major bottleneck is data collection. Teleoperation is particularly challenging for UAMs due to expensive and fragile hardware and unintuitive interface. To address this, recent work explores cross-embodiment collection, with the Universal Manipulation Interface (UMI) [chi2024universal] offering a portable, low-cost way to record demonstrations across environments. By decoupling demonstrations from specific robots, UMI enables training of embodiment-agnostic manipulation policies.
@@
While UMI enables embodiment-agnostic manipulation policies, their success hinges on the embodiment’s ability to execute the generated trajectories. Fixed-base arms with precise controllers are highly “UMI-able [^1] We will formally define “UMI-able” in § IV, and Fig. 6 quantifies how “UMI-able” different embodiments are in simulation.”, able to execute UMI policies as if they were the handheld gripper. In contrast, embodiments like UAMs face stringent physical and control constraints such as stability under aerodynamic disturbances [shi2019neural, o2022neural]. Without accounting for these constraints, UMI policies may yield trajectories that are infeasible, unsafe, or inefficient. Hence, the central challenge is then how to extend UMI beyond highly UMI-able robots to embodiments where control constraints fundamentally shape feasibility.
To address this challenge, we propose Embodiment-Aware Diffusion Policy (EADP), where the key idea is to enable two-way communication between an embodiment-agnostic high-level manipulation policy and embodiment-specific low-level controllers (Fig. LABEL:fig:teaser). Unlike standard UMI systems [chi2024universal, ha2024umilegs] that rely on one-way communication by passing policy outputs directly to controllers, EADP lets low-level controllers actively guide the high-level policy through the denoising process, therefore producing end-effector (EE) trajectories that are more feasible for the target embodiment (e.g., a UAM).
@@
  • We introduce a simulation-based benchmark suite [^2] All code, data, and checkpoints will be publicly released after acceptance. Result videos can be found at umi-on-air.github.io, which facilitates systematic investigation of the embodiment gap when using UMI demonstration data on embodiments with varying UMI-abilities.
@@ II-A Mobile Manipulation
Ground-Based Manipulation. Ground-based mobile manipulation systems traditionally emerged to be designed to tailor to specific use cases [khatib1996force, chestnutt2005footstep, feng2014optimization, krotkov2018darpa, bajracharya2024demonstrating]. This led to a strong reliance on task and motion planning and model-based control, to capture the unique embodiment kinematics and dynamics for the specific mobile system. Recent learning-based systems have shown success in leveraging behavior cloning [fu2024mobile, yang2024equibot, ze2024generalizable, he2024omnih2o, jiang2025behavior], reinforcement learning (RL) [xia2021relmogen, hu2023causal, uppal2024spin, mendonca2024continuously], combining reinforcement learning for locomotion and behavior cloning for manipulation [liu2024visual, ha2024umilegs] as an alternative for ground-based mobile manipulation via imitation learning.
Aerial Manipulation. While aerial manipulation is a subset of mobile manipulation, it introduces distinct challenges compared to ground-based mobile embodiments, including disturbance near ground and wall, stability requirements, underactuated nonlinear dynamics, and strict payload constraints. Aerial manipulation has been demonstrated across a wide variety of applications, including inspection of surfaces [bodie2019omnidirectional, guo2024aerial], writing and painting [lanegger2022aerial, guo2024flying], object grasping [ubellacker2024softgrasp, bauer2024an], insertion [schuster2022automated, wang2023millimeter], and articulated object interaction [brunner2022planning, lee2021manip]. These successes have typically relied on specialized hardware systems coupled with carefully engineered control strategies for a specific task, but remain hard to scale to novel manipulation goals or environments. To support broader deployment across aerial tasks, recent research has shifted toward developing general-purpose frameworks that abstract away embodiment-specific dynamics—for example, EE-centric control interfaces [he2025flying], which decouples high-level decision making from low-level embodiment-specific actuation. Despite these advancements, robust and generalizable policies for such systems require training on large-scale robot data that span diverse objects, scenes, and flight conditions. However, data collection directly with UAMs is challenging due to the difficulty of deploying drones across diverse physical environments, motivating alternative strategies for data collection and deployment.
@@ II-B Cross-embodiment Learning
Recent works have explored using large-scale cross-embodiment datasets as a pretraining strategy, involving data collected from various robotic embodiments and subsequently finetuned to accommodate specific hardware embodiments [liu2024rdt, yang2024pushinglimitscrossembodimentlearning, reed2022a, qiu2025humanoid]. These approaches rely on the assumption of a unified action space, which allows data sharing across robots with similar morphology, primarily robotic arms or mobile manipulators. Despite improvements in generalization, such methods require extensive embodiment-specific finetuning datasets to ensure policies adapt to the target embodiment.
An alternative strategy involves collecting demonstration data directly from the human-embodiment using intuitive handheld interfaces [chi2024universal, shafiullah2023bringing, song2020grasping, young2021visual, xu2025dexumi]. For instance, UMI [chi2024universal] bridges the embodiment gap by minimizing differences between the observation space and action spaces of the hand-held gripper and the robot embodiment. This approach reduces the overhead associated with robotic data collection and enables large-scale, in-the-wild demonstrations without relying on physical robot hardware. While these methods have demonstrated impressive capabilities for tasks including precise or dynamic manipulation, policies trained directly from human demonstrations internalize action constraints reflective of human embodiment. As a result, these policies remain unaware of the distinct dynamics and physical limitations of embodiments like mobile manipulators [ha2024umilegs, he2025flying], where the EE cannot precisely track the generated action sequences, leading to unreliable execution.
To incorporate embodiment information into policy representations, recent works have developed specialized model architectures. Embodiment-aware policies leveraged graph neural networks (GNNs), explicitly modeling robot structures as graphs, with joints as nodes and links as edges [wang2018nervenet, huang2020one]. Subsequent works have explored transformer-based models, motivated by their superior representational capacity [kurin2020my, gupta2022metamorph, sferrazza2024body, patel2024get]. Embodiment-aware architectures have demonstrated impressive zero-shot generalization capabilities within RL contexts by using large-scale training with extensive embodiment randomization, yet their adoption within imitation learning remains limited due to a lack of such data. In contrast, our work proposes incorporating embodiment-awareness during inference time by integrating feedback from a low-level embodiment-specific controller into a diffusion policy’s trajectory generation process. By iteratively guiding diffusion-based sampling toward controller-feasible trajectories, we gain the benefits of abstraction from an ee-centric action space while producing trajectories that respect the robot’s physical constraints.
@@ III-A Data Collection Interface
We adopt the Universal Manipulation Interface (UMI) [chi2024universal] paradigm for human demonstration collection: a lightweight, hand-held gripper with a wrist-mounted camera for egocentric observation, and a shared action interface expressed in the EE frame. This design enables in-the-wild data collection without requiring robot hardware, and aligns the training and deployment modalities by replicating the camera–gripper configuration on the robot.
@@
Each demonstration consists of synchronized egocentric RGB images, 6-DoF EE pose trajectories, and continuous gripper width tracked using fiducial markers on the fingers. These sequences form input–output pairs for policy learning: the input is an observation window consisting of images, relative EE poses, and gripper widths, while the output is a horizon of future actions given as relative EE trajectories and gripper widths. A conditional UNet-based [ronneberger2015u] diffusion policy is trained on these pairs, enabling the generation of multimodal action sequences from the UMI demonstrations.
@@ III-B End-Effector-Centric Controllers
To guide the diffusion policy toward embodiment-feasible behaviors, we define a tracking cost $L_{\text{track}}(a)$ that evaluates how well a given trajectory $a$ can be executed by a particular controller. High tracking cost indicates segments of the trajectory that are hard to follow—due to dynamic infeasibility, underactuation, or control saturation—while low cost indicates better alignment with the embodiment’s capabilities.
@@
Richer controller instantiations can be used for robots that require accurate modeling of dynamics, such as UAMs. We adopt the EE–centric whole-body MPC from [he2025flying]. This controller coordinates UAV and manipulator motion by optimizing a finite-horizon cost function subject to dynamics and actuation constraints. The state and control variables are defined as:
@@
where $\bm{v}\in\mathbb{R}^{6}$ is the body velocity (linear + angular), $\bm{\theta}\in\mathbb{R}^{n}$ are the manipulator’s joint angles, $\bm{\tau}\in\mathbb{R}^{6}$ is the commanded wrench (forces and torques) and $\bm{\theta}_{\text{cmd}}\in\mathbb{R}^{n}$ are the commanded joint angles. Note that we adopt a similar UAM system with [he2025flying], which is a fully-actuated hexarotor, allowing us to send the commanded 6-dim control wrench directly.
@@ III-C Embodiment-Aware Diffusion Guidance
As illustrated in Fig. 3, at inference time, we guide the conditional diffusion policy using this gradient feedback from the low-level controller. Let $a^{k}$ denote the noisy reference trajectory sample at diffusion timestep $k\in\{K,\dots,1\}$, conditioned on observation data $\bm{o}$. We use the standard DDIM [song2020denoising] update step, given by:
@@
We incorporate gradient feedback from the tracking cost $L_{\text{track}}$ defined in Eq. (2, 6) similarly to classifier-based guidance [dhariwal2021diffusion]. Specifically, we apply a guidance step to the trajectory sample toward feasible modes:
@@ IV-B Real-world Experiments
We next address our second research question: does EADP enable reliable transfer of UMI-trained policies to real-world UAMs? To this end, we evaluate on three aerial manipulation tasks (Fig. 2) that span precision, robustness, and long-horizon execution, followed by a cross-environment generalization test. We conducted experiments using a fully actuated hexarotor drone with a $4$ DoF manipulator and a gripper. A motion capture system provides the drone state, while the EE state is computed using forward kinematics. Fig. 8 summarizes results across all trials.
@@ V Conclusion & Discussion
While promising, our work leaves several directions for future research. First, the current system operates with a temporal gap between policy inference (around 1-2 Hz) and high-frequency control (50 Hz). This mismatch could be alleviated through streaming diffusion methods [jiang2025streaming] or continuous guidance mechanisms that allow tighter integration between policy and controller. Second, while we demonstrate EADP with IK and MPC instantiations, the framework is not limited to analytical controllers. It can be naturally extended to learned or reinforcement learning–based controllers using learned dynamics models.
@@ VI Acknowledgment
This work was supported by the Robotics Institute Summer Scholars program. We thank Zeji Yi for his thoughtful discussions, and Yutong Wang and Mohammadreza Mousaei for their help with the experiments. This work was partially funded by NSF Award #2512805, #2143601, #2037101, and #2132519 and Toyota Research Institute. Guanya Shi holds concurrent appointments as an Assistant Professor at Carnegie Mellon University and as an Amazon Scholar. This paper describes work performed at Carnegie Mellon University and is not associated with Amazon.

### v3

@@ Abstract
We introduce UMI-on-Air, a framework for embodiment-aware deployment of embodiment-agnostic manipulation policies. Our approach leverages diverse, unconstrained human demonstrations collected with a handheld gripper (UMI) to train generalizable visuomotor policies. A central challenge in transferring these policies to constrained robotic embodiments—such as aerial manipulators—is the mismatch in control and robot dynamics, which often leads to out-of-distribution behaviors and poor execution. To address this, we propose Embodiment-Aware Diffusion Policy (EADP), which couples a high-level UMI policy with a low-level embodiment-specific controller at inference time. By integrating gradient feedback from the controller’s tracking cost into the diffusion sampling process, our method steers trajectory generation towards dynamically feasible modes tailored to the deployment embodiment. This enables plug-and-play, embodiment-aware trajectory adaptation at test time. We validate our approach on multiple long-horizon and high-precision aerial manipulation tasks, showing improved success rates, efficiency, and robustness under disturbances compared to unguided diffusion baselines. Finally, we demonstrate deployment in previously unseen environments, using UMI demonstrations collected in the wild, highlighting a practical pathway for scaling generalizable manipulation skills across diverse—and even highly constrained—embodiments. All code, data, checkpoints, and result videos can be found at umi-on-air.github.io.
@@ I Introduction
There is a growing interest in extending manipulation beyond the lab and into more complex, dynamic settings. Among emerging embodiments, unmanned aerial manipulators (UAMs) hold particular promise. Essentially a manipulator with practically limitless reach, UAMs can access environments that are otherwise unreachable or unsafe, such as performing infrastructure maintenance atop towers or harvesting crops in cluttered orchards. Multiple research works have focused on those practical tasks, including non-destructive testing [3], painting [39], drilling [9], light bulb installation [16], etc. These applications highlight the potential of UAMs as an embodiment, but scalable visuomotor policy learning for UAMs remains limited.
A major bottleneck is data collection. Teleoperation is particularly challenging for UAMs due to expensive and fragile hardware and unintuitive interface. To address this, recent work explores cross-embodiment collection, with the Universal Manipulation Interface (UMI) [7] offering a portable way to record demonstrations across environments. By decoupling demonstrations from specific robots, UMI enables training of embodiment-agnostic manipulation policies.
@@
While UMI enables embodiment-agnostic manipulation policies, their success hinges on the embodiment’s ability to execute the generated trajectories. Fixed-base arms with precise controllers are highly “UMI-able [^1] We will formally define “UMI-able” in § IV, and Fig. 6 quantifies how “UMI-able” different embodiments are in simulation.”, able to execute UMI policies as if they were the handheld gripper. In contrast, embodiments like UAMs face stringent physical and control constraints such as stability under aerodynamic disturbances [35, 28]. Without accounting for these constraints, UMI policies may yield trajectories that are infeasible, unsafe, or inefficient. Hence, the central challenge is then how to extend UMI beyond highly UMI-able robots to embodiments where control constraints fundamentally shape feasibility.
To address this challenge, we propose Embodiment-Aware Diffusion Policy (EADP), where the key idea is to enable two-way communication between an embodiment-agnostic high-level manipulation policy and embodiment-specific low-level controllers (Fig. LABEL:fig:teaser). Unlike standard UMI systems [7, 15] that rely on one-way communication by passing policy outputs directly to controllers, EADP lets low-level controllers guide the high-level policy through the denoising process, therefore producing end-effector (EE) trajectories that are more feasible for the target embodiment (e.g., a UAM).
@@
  • We introduce a simulation-based benchmark suite [^2] All code, data, checkpoints, and result videos can be found at umi-on-air.github.io, which facilitates systematic investigation of the embodiment gap when using UMI demonstration data on embodiments with varying UMI-abilities.
@@ II-A Mobile Manipulation
Ground-Based Manipulation. Ground-based mobile manipulation systems traditionally emerged to be designed to tailor to specific use cases [6, 10, 21, 1]. This led to a strong reliance on task and motion planning and model-based control, to capture the unique embodiment kinematics and dynamics for the specific mobile system. Recent learning-based systems have shown success in leveraging behavior cloning [11, 46, 49, 17], reinforcement learning (RL) [44, 18, 27], combining reinforcement learning for locomotion and behavior cloning for manipulation [25, 15] as an alternative for ground-based mobile manipulation via imitation learning.
Aerial Manipulation. While aerial manipulation is a subset of mobile manipulation, it introduces distinct challenges compared to ground-based mobile embodiments, including disturbance near ground and wall, stability requirements, underactuated nonlinear dynamics, and strict payload constraints. Aerial manipulation has been demonstrated across a wide variety of applications, including inspection [3, 12], writing and painting [23, 13], object grasping [38, 2], pick-and-place [5], insertion [41], and articulated object interaction [4, 24]. These successes have typically relied on specialized hardware systems coupled with carefully engineered control strategies for specific tasks, but remain hard to scale to novel manipulation goals or environments. Recent research has shifted toward general frameworks that abstract away embodiment-specific dynamics, such as the EE-centric control interface that decouples high-level decision making from low-level embodiment-specific actuation [16]. However, robust and generalizable policies for such systems require large-scale robot data that span diverse objects, scenes, and flight conditions. However, data collection directly with UAMs is challenging due to the difficulty of deploying drones across diverse physical environments, motivating alternative strategies for data collection and deployment.
@@ II-B Cross-embodiment Learning
Recent works have explored large-scale cross-embodiment pretraining, involving data collected from various robotic embodiments and subsequently finetuned to accommodate specific hardware embodiments [26, 47, 31, 30]. These approaches rely on the assumption of a unified action space, which allows data sharing across robots with similar morphology, primarily robotic arms or mobile manipulators. Despite improvements in generalization, such methods require extensive embodiment-specific finetuning datasets to ensure policies adapt to the target embodiment.
An alternative strategy collected demonstrations directly from the human-embodiment using intuitive handheld interfaces [34, 37, 48, 45]. For instance, UMI [7] bridges the embodiment gap by aligning the observation space and action spaces of the hand-held gripper and the robot embodiment, enabling large-scale, in-the-wild demonstrations without relying on physical robot hardware. However, policies trained directly from human demonstrations internalize action constraints reflective of human embodiment. As a result, these policies remain unaware of the distinct dynamics and physical limitations of embodiments like mobile manipulators [15, 16], where the EE cannot precisely track the generated action sequences, leading to unreliable execution.
To incorporate embodiment information into policy representations, recent works developed specialized model architectures. Embodiment-aware policies leveraged graph neural networks (GNNs), explicitly modeling robot structures as graphs, with joints as nodes and links as edges [42, 19]. Subsequent works explored transformer-based models, motivated by their superior representational capacity [22, 14, 33, 29]. Embodiment-aware architectures showed impressive zero-shot generalization capabilities within RL contexts by using large-scale training with extensive embodiment randomization, yet their adoption within imitation learning remains limited due to data scarcity. In contrast, our work proposes incorporating embodiment-awareness during inference time by integrating feedback from a low-level embodiment-specific controller into a diffusion policy’s trajectory generation process. By iteratively guiding diffusion-based sampling toward controller-feasible trajectories, we gain the benefits of abstraction from an ee-centric action space while producing trajectories that respect the robot’s physical constraints.
@@ III-A Data Collection Interface
We adopt the UMI [7] paradigm for human demonstration collection: a lightweight, hand-held gripper with a wrist-mounted camera for egocentric observation, and a shared action interface expressed in the EE frame. This design enables in-the-wild data collection without requiring robot hardware, and aligns the training and deployment modalities by replicating the camera–gripper configuration on the robot.
@@
Each demonstration consists of synchronized egocentric RGB images, 6-DoF EE pose trajectories, and continuous gripper width tracked using fiducial markers on the fingers. These sequences form input–output pairs for policy learning: the input is an observation window consisting of images, relative EE poses, and gripper widths, while the output is a horizon of future actions given as relative EE trajectories and gripper widths. A conditional UNet-based [32] diffusion policy is trained on these pairs, enabling the generation of multimodal action sequences from the UMI demonstrations.
@@ III-B End-Effector-Centric Controllers
To guide the diffusion policy toward embodiment-feasible behaviors, we define a tracking cost $L_{\text{track}}(a)$ that evaluates how well a given trajectory $a$ can be executed by a particular controller. High tracking cost indicates trajectory segments that are difficult to follow due to dynamic infeasibility, underactuation, or control saturation, while low cost reflects better alignment with the embodiment capabilities.
@@
Richer controller instantiations can be used for robots that require accurate modeling of dynamics, such as UAMs. We adopt the EE–centric whole-body MPC from [16], including the formulation and all parameters. The MPC is solved using ACADOS [40]. This controller coordinates UAV and manipulator motion by optimizing a finite-horizon cost function subject to dynamics and actuation constraints. The state and control variables are defined as:
@@
where $\bm{v}\in\mathbb{R}^{6}$ is the body velocity (linear + angular), $\bm{\theta}\in\mathbb{R}^{n}$ are the manipulator’s joint angles, $\bm{\tau}\in\mathbb{R}^{6}$ is the commanded wrench (forces and torques) and $\bm{\theta}_{\text{cmd}}\in\mathbb{R}^{n}$ are the commanded joint angles. Note that we adopt a similar UAM system with [16], which is a fully-actuated hexarotor, allowing us to send the commanded 6-dim control wrench directly.
@@ III-C Embodiment-Aware Diffusion Guidance
As illustrated in Fig. 3, at inference time, we guide the conditional diffusion policy using this gradient feedback from the low-level controller. Let $a^{k}$ denote the noisy reference trajectory sample at diffusion timestep $k\in\{K,\dots,1\}$, conditioned on observation data $\bm{o}$. We use the standard DDIM [36] update step, given by:
@@
We incorporate gradient feedback from the tracking cost $L_{\text{track}}$ defined in Eq. (2, 6) similarly to classifier-based guidance [8]. Specifically, we apply a guidance step to the trajectory sample toward feasible modes:
@@ IV-B Real-world Experiments
We next address our second research question: does EADP enable reliable transfer of UMI-trained policies to real-world UAMs? To this end, we evaluate on three aerial manipulation tasks (Fig. 2) that span precision, robustness, and long-horizon execution, followed by a cross-environment generalization test. We conducted experiments using a fully actuated hexarotor drone with a $4$ DoF manipulator and a gripper. A motion capture system provides the drone state, while the EE state is computed using forward kinematics. Policy inference runs on an NVIDIA RTX 4070Ti at approximately 1.5 Hz; the MPC executes on CPU at 50 Hz, interpolating between asynchronously published reference waypoints. Fig. 8 summarizes results across all trials.
@@ V Conclusion & Discussion
While promising, our work leaves several directions for future research. First, the current system operates with a temporal gap between policy inference (around 1-2 Hz) and high-frequency control (50 Hz). This mismatch could be alleviated through streaming diffusion methods [20] or continuous guidance mechanisms that allow tighter integration between policy and controller. Second, while we demonstrate EADP with IK and MPC instantiations, the framework is not limited to analytical controllers. It can be naturally extended to learned or reinforcement learning–based controllers using learned dynamics models. We note that sampling-based guidance methods such as MPPI [43] offer an alternative to gradient-based steering, and achieve comparable results in simulation but are prohibitively expensive in practice due to repeated MPC evaluations on CPU.
@@ VI Acknowledgment
This work was supported by the Robotics Institute Summer Scholars program. We thank Zeji Yi and Junyi Geng for their thoughtful discussions, and Yutong Wang and Mohammadreza Mousaei for their help with the experiments. This work was partially funded by NSF Award #2512805, #2143601, #2037101, and #2132519 and Toyota Research Institute. Guanya Shi holds concurrent appointments as an Assistant Professor at Carnegie Mellon University and as an Amazon Scholar. This paper describes work performed at Carnegie Mellon University and is not associated with Amazon. Dongjae Lee was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2025-02634317).